OpenStudy (anonymous):

I need help need to finish by 11:00 fan and medal

OpenStudy (anonymous):

what do you need help with?

OpenStudy (anonymous):

Question 1 (Multiple Choice Worth 1 points) (03.02) A sequence has its first term equal to 4, and each term of the sequence is obtained by adding 2 to the previous term. If f(n) represents the nth term of the sequence, which of the following recursive functions best defines this sequence? f(1) = 2 and f(n) = f(n - 1) + 4; n > 1 f(1) = 4 and f(n) = f(n - 1) + 2n; n > 1 f(1) = 2 and f(n) = f(n - 1) + 4n; n > 1 f(1) = 4 and f(n) = f(n - 1) + 2; n > 1 Question 2 (Multiple Choice Worth 1 points) (03.07) The equation of a line is shown below: y = - x - 7 What is the equation of a line which is parallel to this line and passes through (1, 79)? y = -x + 80 y = x - 80 y = 7x - 80 y = -7x + 80 Question 3 (Multiple Choice Worth 1 points) (03.02) The first five terms of a sequence are shown below. 5, 9, 13, 17, 21 If the nth term of this sequence is represented by f(n), which of the following functions best represents this sequence? f(n) = 5n - 1; n ≥ 1 f(n) = 5n + 1; n ≥ 1 f(n) = 4n - 1; n ≥ 1 f(n) = 4n + 1; n ≥ 1 Question 4 (Multiple Choice Worth 1 points) (03.02) If f(x) = 7x - 1, what does f(12) represent? The value if x when (7x - 1) = 12 The value of (7x - 1) when x = 12 The value of 12 f(x) when x = 12 The value of f(x) when (7x - 1) = 12 Question 5 (Multiple Choice Worth 1 points) (03.06) Choose the equation of the horizontal line that passes through the point (2, 3). x = 2 x = 3 y = 2 y = 3 Question 6 (Multiple Choice Worth 1 points) (03.06) A telephone company charges a fixed monthly rate plus a rate per minute of usage. The company charges $155 for 100 minutes of usage and $245 for 200 minutes of usage. An equation can be written to show the relationship between the total minutes used (x) and the total monthly charges (y). Which of the following best describes the steps to draw the graph of y against x? Draw a graph which joins the points (155, 100) and (245, 200) and has a slope = 0.9 Draw a graph which joins the points (155, 100) and (245, 200) and has a slope = 1.1 Draw a graph which joins the points (100, 155) and (200, 245) and has a slope = 1.1 Draw a graph which joins the points (100, 155) and (200, 245) and has a slope = 0.9 Question 7 The description below represents Function A and the table represents Function B: Function A The function is 5 more than 3 times x. Function B x y -1 2 0 5 1 8 Which statement is correct about the slope and y-intercept of the two functions? Their slopes are equal, but y-intercepts are not equal. Their slopes are not equal, but y-intercepts are equal. Both slopes and y-intercepts are equal. Neither slopes nor y-intercepts are equal. Question 8 Maggie has $32 in her money box. She plans to put $9 in her money box every month. If f(n) represents the total amount of money Maggie will have in her money box after n months, which of the following functions represents the relationship between f(n) and n? f(n) = 9n + 32 f(n) = 32n + 9 f(n) = 41n + 9 f(n) = 9n + 41 Question 9 Choose the equation below that represents the line passing through the point (-2, -3) with a slope of -6. y = -6x - 15 y = -6x - 20 y = -6x + 15 y = -6x + 20 Question 10 A student needs to make a rectangular cardboard piece that is 12 inches long. The area of the cardboard piece should be between 60 square inches to 300 square inches. The function f(w) relates the area of the cardboard, in square inches, to the width w in inches. Which of the following shows a reasonable domain for f(w)? 2 ≤ w ≤ 5 5 ≤ w ≤ 25 25 ≤ w ≤ 60 60 ≤ w ≤ 300

OpenStudy (anonymous):

What school is this from?

OpenStudy (anonymous):

Its a home school in London

OpenStudy (texaschic101):

I am not good with function so I will do what I can. 2. y = -x - 7 slope(m) = -1 parallel lines have same slope y = mx + b slope(m) = -1 (1,79)...x = 1 and y = 79 now sub 79 = -1(1) + b 79 = -1 + b 79 + 1 = b 80 = b your equation is : y = -x + 80 ================== 4. f(x) = 7x - 1 The value of 7x - 1 when x is 12. ================== 5. value of horizontal line passing through (2,3) horizontal line has a slope of 0 and is represented by y = a number. so equation is : y = 0x + 3, or just y = 3 ================= 8. f(n) = 9n + 32 ================= 9. (-2,-3)...and slope of -6 y = mx + b -3 = -6(-2) + b -3 = 12 + b -3 - 12 = b -15 = b equation is : y = -6x - 15 ================= I did what I could :)

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

@lxl-rhea-lxl

OpenStudy (anonymous):

1 is b

OpenStudy (anonymous):

3 is either C or D...

OpenStudy (anonymous):

6 is B

OpenStudy (anonymous):

10 is D

OpenStudy (anonymous):

Question 15 (Multiple Choice Worth 1 points) (03.02) Which of the following equations best represents the graph of a function f(m)? f(y) = m f(x) = m y = f(m) x = f(m) Question 16 (Multiple Choice Worth 1 points) (03.05) Arnold deposited some money in a bank. The bank paid him a simple interest every year. After t years, the amount of money f(t), in dollars, that the bank paid Arnold was $1,000 more than 20 times the number of years. Which statement best describes the steps to sketch a graph showing the relationship between f(t) and t? Label t on the x-axis and f(t) on the y-axis, and plot a graph having slope = 20 and y-intercept = 1000. Label t on the x-axis and f(t) on the y-axis, and plot a graph having slope = 1000 and y-intercept = 20. Label f(t) on the x-axis and t on the y-axis, and plot a graph having slope = 20 and y-intercept = 50. Label f(t) on the x-axis and t on the y-axis, and plot a graph having slope = 1000 and y-intercept = 50. Question 17 (Multiple Choice Worth 1 points) (03.05) A video game club charges a fixed annual membership fee of $20 and $2 per video game rented. Let f(n) represent the total annual cost of renting n video games. Which of the following functions best represents the relationship between f(n) and n if the membership was increased by $15 the next year? f(n) = 2n + 35 f(n) = 20n + 2 f(n) = 10n + 35 f(n) = 2n + 10 Question 18 (Multiple Choice Worth 1 points) (03.06) Choose the equation below that represents the line passing through the point (-2, -3) with a slope of -6. y + 3 = -6(x + 2) y + 2 = 6(x + 3) y - 3 = -6(x - 2) y - 2 = 6(x - 3) Question 19 (Fill-In-The-Blank Worth 1 points) (03.05) The graphs of functions f(x) and g(x) = f(x) + k are shown below: graph of line f of x going through ordered pairs 0, 0 and 2, 4. Graph of line g of x going through ordered pairs 0, 2 and 1.5, 5. The value of k is ___. Numerical Answers Expected! Answer for Blank 1: Question 20 (Multiple Choice Worth 1 points) (03.02) If f(x) is the height, in cm, of a sunflower plant which is x days old, which of the following statements best describes the meaning of f(60) = 210? The height of the sunflower plant is 60 cm when it is 210 days old. The height of the sunflower plant is 210 cm when it is 60 days old. The height of the sunflower plant is 210 cm when it is 3.5 days old. The height of the sunflower plant is 60 cm when it is 3.5 days old. Question 21 (Multiple Choice Worth 1 points) (03.01) The tables below show the values of y corresponding to different values of x: Table A x 3 2 1 y 1 0 0 Table B x 3 5 4 y -2 1 1 Which statement is true for the tables? Both Table A and Table B represent functions. Both Table A and Table B do not represent functions. Table A does not represent a function, but Table B represents a function. Table A represents a function, but Table B does not represent a function. Question 22 (Multiple Choice Worth 1 points) (03.07)Choose the equation of the line passing through the point (-3, 1) and parallel to y = -x - 4. y = x + 2 y = -x - 2 y = -x + 4 y = x - 4 Question 23 (Multiple Choice Worth 1 points) (03.07) Choose the equation of the line passing through the point (-2, 4) and perpendicular to y = -negative one halfx + 4. y = 2x + 8 y = 2x y = -2x - 8 y = -2x

OpenStudy (anonymous):

O.O there's more?!?!?!?!?!??!!?!? i hate functions >.<

OpenStudy (anonymous):

I dunno about 15, and 16 is A

OpenStudy (anonymous):

17 is A

OpenStudy (anonymous):

19 is -2

OpenStudy (anonymous):

i dunno 18...

OpenStudy (anonymous):

20 is B

OpenStudy (anonymous):

21 is B

OpenStudy (anonymous):

I dunno 22 or 23...

OpenStudy (anonymous):

Question 24 (Multiple Choice Worth 1 points) (03.05) The graph of a function f(x) is shown below: graph of line going through ordered pairs -2, -1 and 3, 3. An open circle at -2, -1 and a black dot at 3, 3 What is the domain of f(x)? -2 < x ≤ 3 -2 ≤ x < 3 -1 < y ≤ 3 -1 ≤ y < 3 Question 25 (Multiple Choice Worth 1 points) (03.05) Which of the following statements best describes the effect of replacing the graph of f(x) with the graph of f(x) + 4? The graph shifts 4 units up. The graph shifts 4 units down. The graph shifts 4 units left. The graph shifts 4 units right. Question 26 (Multiple Choice Worth 1 points) (03.05) The graph below shows the velocity f(t) of a runner during a certain time interval: graph of line going through ordered pairs 0, 4 and 4, 8. Graph of another line going through ordered pairs 4, 8 and 8, 0. Label on the x axis is time in second and label on the y axis is velocity in meter per second Which of the following describes the intercepts on the graph? The initial velocity of the runner was 4 m/s, and the runner stopped after 8 seconds. The initial velocity of the runner was 8 m/s, and the runner stopped after 4 seconds. The initial acceleration of the runner was 4 m/s2, and the runner stopped after 8 seconds. The initial acceleration of the runner was 8 m/s2, and the runner stopped after 4 seconds. Question 27 (Multiple Choice Worth 1 points) (03.05) The table below shows the distance d(t) in feet that an object travels in t seconds: t (second) d(t) (feet) 3 126 4 224 5 350 6 504 What is the average rate of change of d(t) between 3 seconds and 5 seconds, and what does it represent? 59.5 m/s; it represents the average speed of the object between 3 seconds and 5 seconds 112 m/s; it represents the average speed of the object between 3 seconds and 5 seconds 112 m/s; it represents the average distance traveled by the object between 3 seconds and 5 seconds 59.5 m/s; it represents the average distance traveled by the object between 3 seconds and 5 seconds Question 28 (Multiple Choice Worth 1 points) (03.03) Choose the equation that represents the graph below: Graph of a line passing through points 0 comma 4 and 3 comma 0 y = -negative three fourthsx + 4 y = -negative four thirdsx + 4 y = three fourthsx - 4 y = 4 thirdsx - 4 Question 29 (Multiple Choice Worth 1 points) (03.02) If f(x) = 2x2 + 2, find f(5). 102 52 22 -1.5 Question 30 (Multiple Choice Worth 1 points) (03.03) The location of an ant from a tree was 3 meters west (negative direction). The ant walked towards the tree at a constant rate and reached the tree in 2 minutes. If y represents the location of the ant from the tree, in meters, and x represents the time in minutes, which graph best represents this situation? graph of line going through ordered pairs negative 3, 0 and 0, 2 graph of line going through ordered pairs 0, negative 3 and 2, 0 graph of line going through ordered pairs 0, 2 and 3, 0 graph of line going through ordered pairs 0, negative 2 and negative 3, 0

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

y dont u try asking 1 question at a time

OpenStudy (anonymous):

no one is gonna answer all those questions

OpenStudy (secret-ninja):

Don't lie kid, these questions are from an FLVS exam for 9th grade math. This was a big exam you just cheated on. People have to walk you through it, not give you the answers. This is no "London home-school".

DivineSolar (divinesolar):

Yep i took that not too long ago. @Secret-Ninja

OpenStudy (sparklestaraa):

@CanadianSteve99